OpenStudy (anonymous):

Who is really good at 8th grade math? I will give metals!

OpenStudy (anonymous):

I can help =)

OpenStudy (anonymous):

It depends what topic in 8th grade math.. Im in 8th grade but i take 9th grade math.

OpenStudy (anonymous):

I have 5 questions

OpenStudy (anonymous):

Okay, fire away =)

OpenStudy (anonymous):

give me the questions and ill see if i can

OpenStudy (anonymous):

A set of equations is given below: Equation M: y = 3x + 4 Equation P: y = 3x + 7 Which of the following options is true about the solution to the given set of equations? (4 points) Two solutions Infinite solutions No solution One solution

OpenStudy (anonymous):

Okay so what specifically do you not get? =)

OpenStudy (anonymous):

#4:

OpenStudy (anonymous):

i know some websites that could do it and explain it faster and easier than i could

OpenStudy (anonymous):

A set of equations is given below: Equation S: y = x + 9 Equation T: y = 2x + 1 Which of the following steps can be used to find the solution to the set of equations? (4 points) x + 1 = 2x + 9 x = 2x + 1 x + 9 = 2x x + 9 = 2x + 1

OpenStudy (anonymous):

Okay i can only see problems two and four, just so you know =)

OpenStudy (anonymous):

and five =)

OpenStudy (anonymous):

#1:

OpenStudy (anonymous):

metals? MEDALS lol

OpenStudy (anonymous):

#3:

OpenStudy (anonymous):

Yea Desomos is a really good resource for this type of stuff, but okay so do u know h=the processes to solve these?

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

Nope haha

OpenStudy (anonymous):

I had an appointment when we got taught this

OpenStudy (anonymous):

Question two: Okay so using the substitution method, solve the system of equations =)

OpenStudy (anonymous):

How do I substitute?

OpenStudy (anonymous):

Okay so for the substitution method, what you basically do is you solve for the variable in one of the equations, either one, does not matter, and the u use that value and substitute it for the variable in the other equation. I hope that explains that clear enough =)

OpenStudy (anonymous):

Let me know what you don't get =)

OpenStudy (preetha):

Good work Country Gurl!

OpenStudy (anonymous):

Thanks so much @Preetha =)

OpenStudy (anonymous):

I just got through this lesson so it's kinda helping me solidify it in my head, by helping someone else =)

OpenStudy (anonymous):

I am so confused

OpenStudy (anonymous):

Okay, where are you confused? LOL =)

OpenStudy (anonymous):

This might help = ) http://www.regentsprep.org/regents/math/algebra/AE3/AlgSys.htm

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

How can i help now =)

OpenStudy (anonymous):

im still stuck

OpenStudy (anonymous):

Okay, where are you stuck?

OpenStudy (anonymous):

idk where to start and i need to submit this in 10 min!!!!